OpenStudy (anonymous):

Identify the domain and range and sketch the graph of the function y=-sqrt-x

OpenStudy (anonymous):

\[y=\sqrt{-x}\]

OpenStudy (anonymous):

\[y=-\sqrt{-x}\]

OpenStudy (anonymous):

you can't take the square root of a negative number , solve \[-x\geq 0\] in one step to get the domain

OpenStudy (anonymous):

same domainfor \(y=-\sqrt{-x}\)makes no difference

OpenStudy (anonymous):

as for the range, the square root is greater than or equal to zero, so minus the square root will be less than or equal to zero

OpenStudy (anonymous):

Ok so the domian would be (-infinity, infinity)?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

and no

OpenStudy (misty1212):

do you know how to solve \[-x\geq 0\] for \(x\)?

OpenStudy (anonymous):

No

OpenStudy (misty1212):

if not, that is fine, i will show you

OpenStudy (anonymous):

Ok thanks

OpenStudy (misty1212):

ok change the sign, and flip the inequality

OpenStudy (anonymous):

Ok

OpenStudy (misty1212):

the math teachers call it "multiply by \(-1\)" or sommat, but you just make it positive and change greater than or equal to to less that or equal to

OpenStudy (anonymous):

what @misty1212 said

OpenStudy (anonymous):

don't you divide by -1 and flip the sign?

OpenStudy (misty1212):

lol multiply by -1, divide by -1, potato potahto

OpenStudy (anonymous):

ok so would the domain be -1?

OpenStudy (misty1212):

ooooh kay lets go slow

OpenStudy (misty1212):

the domain is an interval right? not a number

OpenStudy (anonymous):

Yes

OpenStudy (misty1212):

the question is, what interval is it now we know that you cannot take the square root of a negative number, so the input has to be positive right?

OpenStudy (misty1212):

the input in this case is \(-x\) in english we would say that \(-x\) must be greater than or equal to zero in math write \[-x\geq 0\]

OpenStudy (misty1212):

to solve that for \(x\) change the sign so that it is \(x\) instead of \(-x\) and change the inequality as well

OpenStudy (misty1212):

i.e \[-x\geq 0\iff x\leq 0\] and that is your domain

OpenStudy (misty1212):

if you want to write it as an interval that works too, call it \[\huge (-\infty, 0]\]

OpenStudy (anonymous):

Ok so what would the range be?

OpenStudy (misty1212):

ok again we go slow the range is all possible outputs the square root is always greater than or equal to zero you have a minus sign in front, so that means the range will be all numbers less than or equal to zero

OpenStudy (anonymous):

Ok so how would I graph it?

OpenStudy (misty1212):

if you want to write it as an interval, write it the same as the domain in this case they are the same

OpenStudy (misty1212):

how to you graph \(y=-\sqrt{-x}\)?

OpenStudy (misty1212):

is that the question?

OpenStudy (anonymous):

Well I have to sketch the graph of the function

OpenStudy (anonymous):

do you know what \(y=\sqrt{x}\) looks like?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

hmm then you can't really do it maybe look it up

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok be very careful here

OpenStudy (anonymous):

wolfram has given you too much information

OpenStudy (anonymous):

where it says "complex valued plot" change it to "real valued plot"

OpenStudy (anonymous):

Oh ok thanks

OpenStudy (anonymous):

you should see something like that